# Controllable Emotion Generation with Emotion Vectors

### v1

@@ 4.2 Emotion score
We aim to evaluate the effectiveness of emotional vectors (EV) in enhancing the emotional expression of generated sentence through classification models. To achieve this, we employed a Multi-Genre Natural Language Inference (MNLI) model called bart-large-mnli that categorizes each sentence into self-designed classes. Three distinct classes: emotionless, neutral, and emotional are choosen. The primary metric used is the probability assigned to the emotional class on the EQ+ dataset, referred to as the Emotion Probability Score. Details are shown in Appendix C.3. A higher score indicates a greater likelihood that the sentence conveys emotional content.
@@
Table 4: Emotion Probability Scores for different models with $EV^{\text{base}}$ conditioning.
Table 4 presents the Emotion Probability Scores (EPR). The results demonstrate that applying EV conditioning consistently achieves the highest emotion probability across most models. For instance, models such as Llama3.1, Qwen2, and MiniCPM show substantial increases in their Emotion Probability Scores when subjected to 2EV, reaching scores of 1.000, 0.9825, and 0.9950 respectively. Conversely, when EV is reduced to -1EV, the majority of models exhibit a decrease in Emotion Probability Scores, indicating a reduction in emotional intensity. For example, Qwen2 drops from 0.6150 to 0.2550. Similarly, Vicuna’s score decreases from 0.5600 to 0.4075. These findings indicate that emotional vectors (EV) can be leveraged to both enhance and attenuate the emotional content of the model’s output, thereby providing effective control over the emotional expression in the generated text.
Emotion Absolute Score
We next prove that the application of EV not only increases the probability of the model expressing emotions, but also that the application of EVs of different modal lengths will increase the strength of the model expressing emotions. To achieve this goal, we use gpt-4o-mini to give an absolute score of 0-100 for each basic emotion of each output of the model, and design an indicator to represent the absolute strength of the emotion of each output, referred to as the Emotion Absolute Score. The details are shown in the appendix C.4.
@@
Table 5: Emotion Absolute Scores for different models with $EV^{\text{base}}$ conditioning.
Table 5 presents the Emotion Absolute Scores(EAS). The results show that after applying EV, the intensity of emotions expressed by most models has been significantly changed. Even if only 1EV is applied, the EAS of llama3.1, Qwen2.5, Gemma and other models have increased by at least 400%. Even for models with poor EV effects, such as Vicuna, minicpm, etc., the EAS has also increased by about 50%. In contrast, for the case of -1EV, the EAS of llama3.1, Qwen2.5, Gemma and other models have been reduced by nearly 90%, and the EAS of Vicuna, minicpm and other models have also been reduced by about 50%. This shows that the application of EV has a significant impact on The absolute strength of the model’s expressed emotion has a significant effect.
@@ 4.3 Effect of Emotion Vectors
In order to verify the effect of EV on improving the corresponding emotions and the universality of EV in different sizes and different architectures, we selected four models of different sizes in the same series, similar sizes in different architectures, and different sizes in different architectures, as shown in Table 6 as experiments.
We first extracted the EV of the five basic emotions of these models (including anger, disgust, fear, joy, and sadness), and then applied these vectors one, two, and four times on the EQ+ dataset to obtain the output. Then, using the bart classifier mentioned in this article, all sentences were scored in six categories (including the above five basic emotions and the neutral category) using the MultiLabel mode, and finally the probability of each sentence under these six labels was obtained. We averaged the probabilities of the corresponding emotion labels of the sentences under each EV condition of each model to obtain the results in Table 6. For example, for Llama2-7B, the calculation method of its 1EV, anger condition score is that it applies its own 1*anger EV on EQ+ to obtain the result. After passing through the six classifiers, the average of the probabilities of all sentences in the anger category is taken to obtain the corresponding score.
@@
Table 6: Emotion Analysis of Different Models
From the calculation method of these indicators, it can be seen that the larger the indicator is, the better the effect of the model in expressing the corresponding emotion after applying the corresponding EV. From the results in Table 6, it can be seen that the performance of almost all models has been improved by about 1 times at 1EV, and the EV of most models is close to the performance peak at about 2 times. For the special case of minicpm, its EV is close to the performance peak when it is 4 times. Through our own observation, we found that since minicpm’s ability to follow instructions is relatively weak in the stage of extracting EV, the modulus length of the extracted vector is smaller than the activation value of each layer itself, so its ability to affect the output result is weak, so its performance improvement will only increase with the increase of modulus length. For some models, such as Llama2-7B’s performance on fear EV, its performance began to decline at 4EV. After our inspection, we found that this phenomenon is due to the fact that the modulus length of 4EV is too large compared to the modulus length of its own activation value, which excessively affects the decoding process of the model, causing the model to repeat decoding and become a "repeater", thereby affecting the discrimination of the classifier, and then causing the performance to decline.
@@ 4.4 Visualization of Emotion Vectors
Tsne visualization of EV A t-SNE dimensionality reductionVan der Maaten and Hinton (2008) reveals that the Emotion Vectors form distinct clusters, each corresponding to a single task. The t-SNE visualization shown in Fig 3 is generated by concatenating the EVs across all layers, followed by the dimensionality reduction. To provide insights into the individual layers’ contributions, we present the visualizations of single-layer EVs in the appendix C.5 Fig 4. These layer-specific visualizations demonstrate how different layers encode and separate emotional features at varying levels of abstraction.
Variability visualization of EV Fig 5 in the appendix C.5 shows histograms of distances within and across emotion states. It can be seen that vectors within the same emotion are closer than those between different emotions, indicating that our proposed emotion vectors are stable within emotional states and not highly influenced by queries. The vectors are constructed by concatenating vectorss from all layers of the model, reduced to 3 dimensions using t-SNE, and cosine distance is used as the metric.
@@ 5 Conclusion
This paper introduces a novel method for expressing and controlling emotions in large-scale language models (LLMs), addressing a significant gap in emotion control within natural language processing (NLP) tasks. Our approach enables the generation of highly effective and universal emotion vectors via a simple prompting mechanism, without requiring additional training. This allows for the flexible, multi-granular control of emotional outputs. Through extensive experiments, we validate the method’s effectiveness across various LLM architectures and scales, particularly highlighting its superior controllability of diverse emotional expressions. Comparative analysis demonstrates that our method outperforms existing techniques in terms of both emotion accuracy and flexibility.
Limitations
In this paper, we propose a method for controllable emotion generation in LLMs. However, our proposed EmotionQuery dataset only contains 500 entries, which is relatively small. Enlarging the size of the dataset may have better results. Furthermore, we are unable to verify the effectiveness of models larger than 14B due to limited experimental resources and some models with access limitations. Although we experimented with five fundamental emotions, we believe that a broader range of emotions, as well as capabilities related to role-playing, can be incorporated into the model using this approach. However, due to limitations in time and resources, we were unable to extend our experiments to include these additional aspects.
Acknowledgments
This work was supported by the Key R&D Program of Zhejiang (2024C01036). The authors would like to thank Ziyue Wang for the encouragement and support while the project. We would also like to thank YunQing Gong, Meixin Liu and Zhiqi Zheng for their helpful comments and discussions on the project.
@@ Appendix A Model Name
Abbreviation Full Name
Llama3.1 Meta-Llama-3.1-8B-Instruct
Llama2 Llama-2-7b-chat-ms
Llama2-13B Llama-2-13b-chat-ms [^2] https://www.modelscope.cn/models/modelscope/Llama-2-13b-chat-ms
Qwen2.5 Qwen2.5-7B-Instruct
Qwen2.5-7B Qwen2.5-7B-Instruct
Qwen2 Qwen2-7B-Instruct
Qwen1.5 Qwen1.5-7B-Chat
baichuan2 Baichuan2-7B-Chat
Yi Yi-6B-Chat
Vicuna vicuna-7b-v1.5
Gemma gemma-7b
MiniCPM MiniCPM3-4B
@@ B.1 EmotionQuery Dataset
The **EmotionQuery** dataset consists of 500 unique queries, distributed across five emotional states: **joy**, **anger**, **disgust**, **fear**, and **sadness**. These emotions are derived from Ekman’s model of basic emotionsEkman (1992), and they serve as the foundational emotional responses for the dataset. For each emotional state $e_{k}$, 100 queries were generated, resulting in a total of 500 queries.
@@
The queries were generated using the GPT-4O-mini model OpenAI (2024) through the following process:
@@ B.2 EmotionQuery+ Dataset
  • 150 additional queries were generated using the GPT-4O-mini model OpenAI (2024) with a new prompt designed to elicit neutral, everyday communication. These queries are not intended to provoke any emotional response, but rather represent common, neutral questions or statements encountered in daily life.
@@
"Can you provide the details in writing?",
"How do you ensure quality in your work?",
"Is there a form I need to fill out?",
"What are the safety procedures here?",
"How do we track our progress?"

### v2

@@ 5.2 Emotion score
We aim to evaluate the effectiveness of emotional vectors (EV) in enhancing the emotional expression of generated sentence through classification models. To achieve this, we employed a Multi-Genre Natural Language Inference (MNLI) model called bart-large-mnli that categorizes each sentence into self-designed classes. Three distinct classes: emotionless, neutral, and emotional are choosen. The primary metric used is the probability assigned to the emotional class on the EQ+ dataset, referred to as the Emotion Probability Score. Details are shown in Appendix D.3. A higher score indicates a greater likelihood that the sentence conveys emotional content. Table 3(a) presents the Emotion Probability Scores (EPR). The results demonstrate that applying EV conditioning consistently achieves the highest emotion probability across most models. For instance, models such as Llama3.1, Qwen2, and MiniCPM show substantial increases in their Emotion Probability Scores when subjected to 2EV, reaching scores of 1.000, 0.9825, and 0.9950 respectively. Conversely, when EV is reduced to -1EV, the majority of models exhibit a decrease in Emotion Probability Scores, indicating a reduction in emotional intensity.
Emotion Absolute Score
We next prove that the application of EV not only increases the probability of the model expressing emotions, but also that the application of EVs of different modal lengths will increase the strength of the model expressing emotions. To achieve this goal, we use gpt-4o-mini to give an absolute score of 0-100 for each basic emotion of each output of the model, and design an indicator to represent the absolute strength of the emotion of each output, referred to as the Emotion Absolute Score. The details are shown in the appendix D.4.
@@
((a)) Emotion Probability Scores with $EV^{\text{base}}$.
@@
((b)) Emotion Absolute Scores with $EV^{\text{base}}$.
Table 3: Comparison of Emotion Probability and Absolute Scores across models.
Table 3(b) presents the Emotion Absolute Scores(EAS). The results show that after applying EV, the intensity of emotions expressed by most models has been significantly changed. Even if only 1EV is applied, the EAS of llama3.1, Qwen2.5, Gemma and other models have increased by at least 400%. In contrast, for the case of -1EV, the EAS of llama3.1, Qwen2.5, Gemma and other models have been reduced by nearly 90%.
@@ 5.3 Effect of Emotion Vectors
To evaluate the effectiveness and generalizability of Emotion Vectors (EVs) across different model architectures and sizes, we conduct a comparative study on four representative models. These models were selected to cover: (1) different sizes within the same architecture family, (2) similar sizes across different architectures, and (3) diverse sizes and architectures. Details are shown in Table 4.
For each model, we extracted EVs corresponding to five basic emotions (anger, disgust, fear, joy, and sadness), and applied them at different intensities (1×, 2×, and 4×) on the EQ+ dataset. To quantify emotional expression under different EV settings, we introduce the Target Emotion Confidence (TEC) score, which measures how confidently a classifier identifies the intended emotion in the generated response. A higher TEC score indicates better alignment with the target emotion after EV application. The results are summarized in Table 4.
Target Emotion Confidence $\uparrow$
@@
Table 4: Target Emotion Confidence (TEC, $\uparrow$ better) scores of different models on five basic emotions. For each model, we apply Emotion Vectors (EVs) corresponding to each emotion at varying intensities (0×, 1×, 2×, 4×) on the EQ+ dataset.
From Table 4, we observe that for most models, applying 1× or 2× EV significantly enhances the emotional alignment, with diminishing returns or even slight degradation at 4× intensity. For instance, LLaMA2-7B achieves strong improvements at 1× and 2× EV, but experiences a drop under 4× fear EV. Upon inspection, this is due to excessively large EV magnitude relative to the model’s activation scale, which interferes with decoding and leads to repetitive outputs that confuse the classifier.
A detailed explanation of the TEC computation process can be found in Appendix D.5.1.
5.4 Controllability Under Emotionally Biased Prompts
To further evaluate the robustness and precision of our emotion control method, we separately re-calculate the TEC score of Qwen-2.5 on EQ+ dataset where the input prompts themselves carry strong emotional tendencies. Such prompts naturally bias the model’s generation toward a particular emotion. The goal is to assess whether our Emotion Vectors (EVs) can override this inherent bias and reliably guide the output toward a specified target emotion.
[…]
Quantitative Evaluation
We compile 5 tables, one for each target emotion, where:
  • Rows indicate the original emotion of the input query (from EQ+);
[…]
  • Cell values denote the average classifier confidence for the target emotion.
Figure 2 shows an example matrix for the target emotion Anger. As EV intensity increases, the model consistently produces outputs that better align with the target emotion—even when the prompt is biased toward a different emotion.
Figure 2: Target Emotion Confidence (TEC) scores across different Emotion Vector (EV) intensities for each target emotion. Each subplot corresponds to a specific target emotion (e.g., anger, joy), and each line represents the TEC score achieved when applying the EV to prompts originally associated with a given emotion.
The full set of emotion-specific matrices is provided in Appendix D.5.2.
@@ 5.5 Visualization of Emotion Vectors
Tsne visualization of EV A t-SNE dimensionality reduction[58] reveals that the Emotion Vectors form distinct clusters, each corresponding to a single task. The t-SNE visualization shown in Fig 5 is generated by concatenating the EVs across all layers, followed by the dimensionality reduction. To provide insights into the individual layers’ contributions, we present the visualizations of single-layer EVs in the Fig 3. These layer-specific visualizations demonstrate how different layers encode and separate emotional features at varying levels of abstraction.
Variability visualization of EV Fig 4 shows histograms of distances within and across emotion states. It can be seen that vectors within the same emotion are closer than those between different emotions, indicating that our proposed emotion vectors are stable within emotional states and not highly influenced by queries. The vectors are constructed by concatenating vectorss from all layers of the model, reduced to 3 dimensions using t-SNE, and cosine distance is used as the metric.
@@ 6 Conclusion
While emotion serves as a foundational element in human-centered domains such as education, healthcare, and mental health, current LLMs remain fundamentally limited in their capacity to generate emotionally attuned and contextually appropriate responses. These limitations hinder their effectiveness as truly collaborative agents in affect-sensitive settings. To address challenges, we propose a novel and universally applicable approach that enables fine-grained emotional control in LLMs through the use of an Emotion Vector. By extracting this intrinsic emotional representation via simple prompting and integrating it during inference, our method allows for precise and controllable emotional expression without compromising textual quality. Extensive evaluations across multiple model architectures confirm that the approach yields consistent, stable, and context-aware emotional outputs, overcoming the model-specific constraints that have hampered prior efforts. This work not only provides a scalable and efficient technical pathway toward emotionally intelligent agents but also offers deeper insights into how such models can be endowed with robust affective capabilities, paving the way for more meaningful and effective human–AI interactions in emotionally consequential domains.
[…]
Acknowledgements
This work was supported by the Key R&D Program of Zhejiang (2024C01036). The authors would like to express their sincere gratitude to Ziyue Wang for the continuous encouragement and support throughout the project, and to Jirui Dai for the valuable assistance in revising the figures in the article. We would also like to thank Yunqing Gong, Meixin Liu, and Zhiqi Zheng for their insightful comments and helpful discussions that greatly contributed to this work.
Declaration
Author Contributions. The authors declare the following contributions to this work: Yurui Dong and Luozhijie Jin conceived the main idea and conducted all the experiments. Zhi Liu and Jiaxi Yang provided guidance and valuable advice throughout the work. Yao Yang and Bingjie Lu are affiliated with the Key R&D Program of Zhejiang (2024C0103), which supported this study.All authors provided critical feedback and helped shape the research, analysis and manuscript. All authors discussed the results and contributed to the final manuscript.
The datasets and code we developed in this work can be found at the open accessed github project as folloing: https://github.com/xuanfengzu/EmotionVector.
@@ Appendix A Model Name
The model name and references are shown in table 5.
Abbreviation Full Name Reference
Llama3.1 Meta-Llama-3.1-8B-Instruct Dubey et al. [57]
Llama2 Llama-2-7b-chat-ms Touvron et al. [59]
Llama2-13B Llama-2-13b-chat-ms11footnotemark: 1 Touvron et al. [59]
Qwen2.5 Qwen2.5-7B-Instruct Yang et al. [60]
Qwen2 Qwen2-7B-Instruct Yang et al. [61]
Qwen1.5 Qwen1.5-7B-Chat Bai et al. [62]
Qwen1 Qwen-7B-Chat Bai et al. [62]
baichuan2 Baichuan2-7B-Chat Yang et al. [63]
Yi Yi-6B-Chat Young et al. [64]
Vicuna vicuna-7b-v1.5 Chiang et al. [65]
Gemma gemma-7b Team et al. [66]
MiniCPM MiniCPM3-4B Hu et al. [67]
@@
 [^2] https://www.modelscope.cn/models/modelscope/Llama-2-13b-chat-ms
@@ B.1 EmotionQuery Dataset
The **EmotionQuery** dataset consists of 500 unique queries, distributed across five emotional states: **joy**, **anger**, **disgust**, **fear**, and **sadness**. These emotions are derived from Ekman’s model of basic emotions[55], and they serve as the foundational emotional responses for the dataset. For each emotional state $e_{k}$, 100 queries were generated, resulting in a total of 500 queries.
@@
The queries were generated using the GPT-4O-mini model [56] through the following process:
@@ B.2 EmotionQuery+ Dataset
  • 150 additional queries were generated using the GPT-4O-mini model [56] with a new prompt designed to elicit neutral, everyday communication. These queries are not intended to provoke any emotional response, but rather represent common, neutral questions or statements encountered in daily life.
@@
”Can you provide the details in writing?”,
”How do you ensure quality in your work?”,
”Is there a form I need to fill out?”,
”What are the safety procedures here?”,
”How do we track our progress?”